# Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v3

@@ Abstract
(This is the second version of this working paper. A more detailed follow-up with interesting applications is in preparation.) We develop a family of self-normalized concentration inequalities for marginal mean under martingale-difference structure and $\phi/\tilde{\phi}-$mixing conditions, where the latter includes many processes that are not strongly mixing. The variance term is fully data-observable: naive sample variance in the martingale case and an empirical block long-run variance under mixing conditions. Thus, no predictable variance proxy is required. No specific assumption on the decay of the mixing coefficients (e.g. summability) is needed for the validity. The constants are explicit and the bounds are ready to use.
@@ 1.1 Self-Normalized Inequality
They also proved that (3) holds for “heavy-on-left” MDS as well. Compared with (1), (2)-(4) hold under much weaker moment conditions but their statistical applications are not so convenient as (1). Meanwhile, the dependence on $\mathbf{M}]_{n}$ appears inside the Laplace transform or in the event $\{[\mathbf{M}]_{n}\leq y\}$, and there is no closed-form expression of the type $\sqrt{[\mathbf{M}]_{n}\log(\frac{1}{\delta})}$ with an explicit confidence parameter $\delta$. However, the variance ($\sigma$) and boundedness ($b$) are usually unknown to us, which reject the direct application of (1).
Under some circumstances, the value of boundedness can still be known from prior information. But, compared with boundedness, it is more challenging and risky to assume the variance to be known. Thus, it would be very ideal to derive an inequality having form (1) but the $\sigma$ is observable (or data-dependent), which is actually the core task of a parallel but tightly connected area, empirical Bernstein inequality of marginal mean.
@@ 1.2 Empirical Bernstein Inequality of Marginal Mean
Based on a variance estimator, ? (?) derived the following EB inequality for IID random variables taking values in $[0,1]$. However, compared with (1), their inequality is not sharp. i.e. The length of one sided confidence interval, denoted as $L_{n}$, does NOT satisfies
@@
where $(V_{i})_{i\geq 1}$ is a variance process adapted to filter $\{\sigma(Z_{1},...,Z_{i})\}_{i\geq 1}$. Function $u$ is essentially the sub-exponential uniform boundary and is user-defined (see Sections 2 and 3 of ? (?) for more details). It is obvious that, when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]$ is invariant with respect to $i$ (e.g. IID), (6) becomes an anytime EB inequality of marginal mean. But it has to pay for a iterated logarithmic loss of the sharpness since it is dedicated to be stopping-time valid. Thus, it is still open whether we can get a sharp EB inequality of marginal mean even for IID data. More recently, ? (?) (Theorem 2, Remark 1) gives a nice answer to this question (Their results include more general weighted empirical mean but we here only focus on simple empirical mean for the ease of comparison). By assuming $Z_{i}$’s as bounded random variables taking value in $[0,1]$, they proved the following EB inequality of marginal mean for all $\delta>0$ when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]=\mu$,
@@
In summary, the results of Ramdas and co-authors are indeed elegant. But, as we discussed for fixed-n EB inequality of marginal mean, they have three structural limitations: (i) they require a known and tight bound on the support length to calculate the confidence interval; (ii) they primarily target the predictable mean (the arithmetic average of conditional expectations), and in general mixing settings there is no clear high-probability control of the gap between this predictable mean and the marginal mean—unless one assumes that all conditional expectations are constant, which essentially reduces to the martingale-difference case. Their recent important work, e.g. ? (?) and ? (?), still rely on the assumption of constant conditional expectation. (iii) they rely on a carefully tuned variance proxy satisfying the following three points simultaneously, (i) predictable; (ii) ensure the super-martingale property; (iii) converge to the true variance. The design of this proxy is highly non-trivial even for martingale differences or i.i.d. data. For some complex mixing processes, whose relevant variance of partial sum is the long-run variance rather than the marginal variance, the difficulty becomes even more significant.
@@ 1.3 Our Contributions
    This inequality matches the classical Bernstein–Freedman constants while allowing the random predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ to appear directly in the bound. It provides a one-parameter form that does not follow from the traditional two-parameter or ratio-type self-normalized inequalities. We then upgrade this to a more useful empirical Bernstein inequality by replacing $\langle\mathbf{M}\rangle_{n}$ with $\mathbf{M}]_{n}$ (see (9)). The empirical version enlarges the constant only slightly, yet keeps the variance and rate dependence optimal, and crucially requires no variance proxy. These inequalities serve as the probabilistic backbone of all subsequent results, allowing us to derive empirical Bernstein bounds for much more complex dependent processes (see C2–C4).
@@
    (Concentration for the marginal mean under complex dependence) We focus on concentration for the marginal mean in the presence of temporal dependence, rather than on the predictable (drifting) mean. In martingale-difference or i.i.d. settings these notions coincide, but in general mixing models, the gap between the two different kinds of means is hard to control. Our inequalities therefore are complement of the whole aforementioned important research of empirical Bernstein inequalities.
@@ 2.1 Martingale Difference Sequence (MDS)
In this subsection we first develop our self-normalized inequalities in the martingale-difference setting. As discussed in Section 1, most existing self-normalized concentration inequalities for the marginal mean either rely on the rarely observable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$,or are formulated for ratio-type statistics such as $\mathbf{M}_{n}/[\mathbf{M}]_{n}$. To the best of our knowledge, the following Theorem 1 provides the first sharp (in the sense of (5)) self-normalized concentration inequality for the marginal mean of a martingale difference sequence (MDS), with a fully observable variance term, which makes it also an empirical Bernstein inequality of MDS.
@@
It is still unclear whether the results in the previous research of self-normalized inequality could imply our (8) as a natural corollary without non-trivial work. Except for the literature mentioned in Section 1, ? (?) discuss self-normalized concentration phenomenon by only requiring $\mathbb{E}[\exp(\lambda A-\frac{1}{2}\lambda^{2}B^{2})]\leq 1$. But ? only concentrated on delivering moment or probabilistic inequalities of ratio-type statistics, which is hard to be extended to our settings directly. More recently, ? (?) talked about two parameter self-normalized inequalities. More specifically, by denoting $S_{n}(a)=c(a)\langle\mathbf{M}\rangle_{n}+[\mathbf{M}]_{n}$ ($c(a)$ a constant depending solely on $a>1/8$), ? gave a series exponential inequalities for ratio statistics (i.e. $\frac{\mathbf{M}_{n}}{S_{n}(a)}$) and two-parameter situations (e.g. ? (?). Hence, under martingale-difference assumption, it seems that our (8) is the first “one-parameter” (only a $t>0$ is needed) exponential inequality whose constants matche with classical Bernstein or Freedman inequality.
As we can see in (9), combines the Bernstein–type tradeoff of Freedman’s inequality with a fully empirical self-normalization. It replaces the unobservable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ by the quadratic variation $\mathbf{M}]_{n}$, which is completely determined by the data. Generally speaking, Theorem 1 will serve as the basic building block for our extensions to general mixing settings (Section 2.2), where $\mathbf{M}]_{n}$ is replaced by a block-based empirical long-run variance. A more acceptable setting similar to MDS is that $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ holds for every $i$, which makes $(Z_{i}-\mu,\mathcal{A}_{i})$ an MDS. However, when $\mu\neq 0$, $\mathbf{M}]_{n}$ becomes a term containing unknown parameter. Using empirical mean to replace $\mu$ yields the following empirical Bernstein inequality.
@@ 3 Conclusions
This paper develops a simple route to empirical Bernstein–type concentration for dependent data. For bounded martingale difference sequences, we obtain a one–parameter Freedman–style inequality with predictable quadratic variation, and then show how to pass from the predictable quadratic variation to the empirical quadratic variation. This yields a fully empirical self–normalized bound in which the variance term is entirely data–driven. For $\phi/\tilde{\phi}-$mixing sequences, we combine a blocking scheme with the martingale results to derive analogous empirical Bernstein inequalities based on block empirical long–run variance, without imposing explicit decay rates on the mixing coefficients.
A technical ingredient underlying our approach is a conditional measurability mechanism, formalized in Lemmas 3 and 4, which allows us to rigorously justify the use of random, data–dependent tilting parameters in exponential supermartingale arguments. This mechanism is independent of any particular model and may be of separate interest for other self–normalized or “empirical variance” inequalities under dependence.
The present paper focuses on the core probabilistic tools and their basic forms. In ongoing work, we apply these inequalities to concrete statistical procedures—such as adaptive inference and sequential methods under temporal dependence—showing that the empirical normalization leads to practical, fully data–driven uncertainty quantification without prior knowledge of long–run variance.
@@ 4 Acknowledgement
The author is grateful to Prof. Dr. Aaditya Ramdas for pointing out an important mistake in the earlier version of this work.

### v4

# Self-Normalized Concentration Inequality For Dependent Data With Sample Variance Only

@@ Abstract
(This is the third version of this working paper.) We develop a family of self-normalized concentration inequalities for marginal mean under martingale-difference structure and $\phi/\tilde{\phi}-$mixing conditions, where the latter includes many processes that are not strongly mixing. The variance term is fully data-observable: naive sample variance in the martingale case and an empirical block long-run variance under mixing conditions. Thus, no predictable variance proxy is required. No specific assumption on the decay of the mixing coefficients (e.g. summability) is needed for the validity. The constants are explicit and the bounds are ready to use.
@@ 1.1 Self-Normalized Inequality
They also proved that (3) holds for “heavy-on-left” MDS as well. Compared with (1), (2)-(4) hold under much weaker moment conditions but their statistical applications are not so convenient as (1). Meanwhile, the dependence on $\mathbf{M}]_{n}$ appears inside the Laplace transform or in the event $\{[\mathbf{M}]_{n}\leq y\}$, and there is no closed-form expression of the type $\sqrt{[\mathbf{M}]_{n}\log(\frac{1}{\delta})}$ with an explicit confidence parameter $\delta$. However, the variance ($\sigma$) and boundedness ($b$) are usually unknown to us, which reject the direct application of (1). More recently, ? (?) did a magnificent work by delivering a class of exponential bounds for the probability that a martingale sequence crosses a time-dependent linear threshold. Generally speaking, their work strengthens nearly all of the aforementioned results to a sharper level under more general conditions. For one-parameter self-normalized inequalities involving a martingale $\mathbf{M}_{n}$ and $\langle\mathbf{M}\rangle_{n}$, the key contribution of ?’s paper is to give a single time-uniform (infinite-horizon) Chernoff/line-crossing framework that systematically produces bounds where deviation of $\mathbf{M}_{n}$ is controlled by a variance processes $V_{n}$ built from $\langle\mathbf{M}\rangle_{n}$. (More specific connection between ? and our paper would be discussed in the incoming sections.)
It is obvious that, under some circumstances, the value of boundedness can still be known from prior information. But, compared with boundedness, it is more challenging and risky to assume the variance to be known. Thus, it would be very ideal to derive an inequality having form (1) but the $\sigma$ is observable (or data-dependent), which is actually the core task of a parallel but tightly connected area, empirical Bernstein inequality of marginal mean.
@@ 1.2 Empirical Bernstein Inequality of Marginal Mean
Based on a variance estimator, ? (?) derived the following EB inequality for IID random variables taking values in $[0,1]$. However, compared with (1), their inequality is not sharp. i.e. The length of one sided confidence interval, denoted as $L_{n}$, does NOT satisfy
@@
where $(V_{i})_{i\geq 1}$ is a variance process adapted to filter $\{\sigma(Z_{1},...,Z_{i})\}_{i\geq 1}$. Function $u$ is essentially the sub-exponential uniform boundary and is user-defined (see Sections 2 and 3 of ? (?) for more details). It is obvious that, when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]$ is invariant with respect to $i$ (e.g. IID), (6) becomes an anytime EB inequality of marginal mean. Because (6) is designed to be stopping-time valid, its fixed-time instantiation typically incurs an iterated-logarithm (or other uniform-boundary) overhead relative to sharp fixed-$n$ bound. Consequently, obtaining sharp fixed-$n$ empirical Bernstein bounds for the marginal mean is nontrivial; a sharp resolution in the bounded case was later provided by ? (?). More specifically, Theorem 2, Remark 1 of ? gives a nice answer to this question (Their results include more general weighted empirical mean but we here only focus on simple empirical mean for the ease of comparison). By assuming $Z_{i}$’s as bounded random variables taking value in $[0,1]$, they proved the following EB inequality of marginal mean for all $\delta>0$ when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]=\mu$,
@@
In summary, the results of Ramdas and co-authors are indeed elegant. But, as we discussed for fixed-n EB inequality of marginal mean, they primarily target the predictable mean (the arithmetic average of conditional expectations). Unfortunately, in general mixing settings, there is no clear high-probability control of the gap between this predictable mean and the marginal mean. Unless one assumes that all conditional expectations are constant, which essentially reduces to the martingale-difference case. Their recent important work, e.g. ? (?) and ? (?), still rely on the assumption of constant conditional expectation. More specifically, it seems that their sharpness rely on a carefully tuned variance proxy satisfying the following three points simultaneously, (i) predictable; (ii) ensure the super-martingale property; (iii) converge to the true variance. Thus, the design of this proxy is non-trivial even for martingale differences or i.i.d. data. For some complex mixing processes, whose relevant variance of partial sum is the long-run variance rather than the marginal variance, it is highly possible the difficulty becomes more significant.
@@ 1.3 Our Contributions
    This inequality can be regarded as 1-sub-Gamma with variance process $(\langle M\rangle_{t})_{t\geq 1}$ and scale parameter $c=b/3$ in the sense of ? (?). But our contribution here is that we give a more elementary but rigorous proof, which relies on the idea of randomizing the parameter $\lambda$ in exponential super-martingale. The empirical version enlarges the constant only slightly, yet keeps the variance and rate dependence optimal, and crucially requires no variance proxy. These inequalities serve as the probabilistic backbone of all subsequent results, allowing us to derive empirical Bernstein bounds for much more complex dependent processes (see C2–C4).
@@
    (Concentration for the marginal mean under complex dependence) We focus on concentration for the marginal mean in the presence of temporal dependence, rather than on the predictable (drifting) mean. In martingale-difference or i.i.d. settings these notions coincide, but in general mixing models, the gap between the two different kinds of means is hard to control. Our inequalities therefore are complementary to the whole aforementioned important research of empirical Bernstein inequalities.
@@ 2.1 Martingale Difference Sequence (MDS)
In this subsection, we first develop the following self-normalized inequalities in the martingale-difference setting.
@@
In the martingale-difference setting with bounded increments $|Z_{i}|\leq b$, our process $(\mathbf{M}_{t},\langle\mathbf{M}\rangle_{t})_{t\geq 1}$ is 1-sub-Gamma with variance process $(\langle M\rangle_{t})_{t\geq 1}$ and scale parameter $c=b/3$ in the sense of ? (?). Specializing their general sub-gamma Chernoff bounds (Theorem 1 and Corollary 1(c)) to a single time point $t=n$ and this particular sub-gamma process yields the one-sided version of (8). Thus, we do not claim (8) as a new martingale inequality per se. Our contribution here is rather a very short, self-contained proof of (8) based on a randomized-parameter mgf argument.
As we can see, (9) combines the Bernstein–type tradeoff of Freedman’s inequality with a fully empirical self-normalization. It replaces the unobservable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ by the quadratic variation $\mathbf{M}]_{n}$, which is completely determined by the data. Generally speaking, Theorem 1 will serve as the basic building block for our extensions to general mixing settings (Section 2.2), where $\mathbf{M}]_{n}$ is replaced by a block-based empirical long-run variance. A more acceptable setting similar to MDS is that $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ holds for every $i$, which makes $(Z_{i}-\mu,\mathcal{A}_{i})$ an MDS. However, when $\mu\neq 0$, $\mathbf{M}]_{n}$ becomes a term containing unknown parameter. Using empirical mean to replace $\mu$ yields the following empirical Bernstein inequality.
@@ 2.2.2 Inequalities
Remark 2
(Modularity and plug-in improvements.) Our dependent-data results are deliberately modular: after the blocking/martingale reduction, all subsequent bounds rely on a single martingale-difference empirical Bernstein (MDS-EB) inequality. Consequently, any sharper or time-uniform MDS-EB tool developed in the future can be plugged into our pipeline with minimal changes; such substitutions primarily refine the numerical constants in the resulting radii while preserving the overall structure (data-driven variance normalization, blocking, and dependence handling). We also emphasize that “sharper” at the level of a standalone MDS inequality does not necessarily imply uniformly shorter confidence intervals once translated to a concrete estimator, since the translation may impose estimator-specific constraints (e.g., fixed weight/normalization, admissible tuning choices, or additional remainder terms). We therefore view improved MDS-EB tools as complementary drop-in components rather than universally dominating replacements.
Remark 3
Many estimators considered can be written as normalized weighted sums of centered observations, i.e.,
$\frac{\sum_{i=1}^{n}w_{i}Z_{i}}{\sum_{i=1}^{n}w_{i}},\ \mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=0,$
where $(Z_{i},\mathcal{A}_{i})$ is the martingale difference sequence defined in Theorem 1. Weight $(w_{i})_{i\geq 1}$ is deterministic or predictable with respect to filteration $(\mathcal{A})_{i\geq 0}$. They may be allowed to be zero since we can impose restriction $\{i:w_{i}>0\}$ to the index set without changing the normalized weighted average. Our dependence-handling arguments (blocking/martingale reduction) are modular in the sense that they reduce the main probabilistic step to a martingale-difference empirical Bernstein (MDS-EB) bound for such weighted sums. Consequently, one may replace the MDS-EB component by alternative tools tailored to the weighted structure, and obtain the corresponding dependent-data inequalities with nearly the same proof. In particular, for weighted averages it is natural to plug in sharp, weight-aware MDS-EB inequalities such as Theorem 4.2 of ? (?). Doing so immediately yields a mixing analogue for normalized weighted sums with sample-variance-only EB bound. We emphasize that this substitution affects only the MDS-EB ingredient while the rest of the construction and the dependence extension remain unchanged.
@@ 3 Conclusions
This paper develops a simple route to empirical Bernstein–type concentration for dependent data. For bounded martingale difference sequences, we give a simple method to obtain one–parameter Freedman–style inequality with predictable quadratic variation, and then show how to pass from the predictable quadratic variation to the naive (blocked) sample variance. This yields a fully empirical self–normalized bound in which the variance term is entirely data–driven. For $\phi/\tilde{\phi}-$mixing sequences, we combine a blocking scheme with the martingale results to derive analogous empirical Bernstein inequalities based on block empirical long–run variance, without imposing explicit decay rates on the mixing coefficients.
@@ 4 Acknowledgement
The author is grateful to Prof. Dr. Aaditya Ramdas for pointing out an important mistake in the earlier version of this work. We also thank him for the guide of literature review.